OpenStudy (yumyum247):

i need a math wizard for help.. am looking for a polynomial equation in standard form from a given table of values..

OpenStudy (yumyum247):

http://imgur.com/zEFVfRO

OpenStudy (yumyum247):

http://imgur.com/jBYO00N

satellite73 (satellite73):

you got the 2 and the -24 already right ?

OpenStudy (yumyum247):

yes

satellite73 (satellite73):

shouldn't be hard to find \(b\) and \(c\) really only need two more pieces of information

OpenStudy (yumyum247):

am stuck on the second picture

satellite73 (satellite73):

me too put \(x=-1\) in \[2x^3+bx^2+cx-24\] set it equal to \(-8\)

OpenStudy (yumyum247):

i did that but that process is toooooooo long...X(

satellite73 (satellite73):

you get \[-2+b-c-24=-8\] or \[b-c=18\]

satellite73 (satellite73):

not as long as the other method

OpenStudy (yumyum247):

ok finish it off plz... lol

OpenStudy (rootbeer003):

hes busy talking to me

OpenStudy (yumyum247):

so what do i do after that

satellite73 (satellite73):

put \(x=1\) set the result equal to -36

satellite73 (satellite73):

\[2+b+c-24=-36\] or \[b+c=-34\]

satellite73 (satellite73):

ok that was a mistake

OpenStudy (yumyum247):

ok

satellite73 (satellite73):

\[2+b+c-24=-36\] isi correct

OpenStudy (yumyum247):